OpenStudy (anonymous):

Trying to find the equation of a parabola. Got it to x^2=-6y+2 but now I don't know what to do with the 2. Help?

OpenStudy (anonymous):

subtract 2 from both sides?

OpenStudy (anonymous):

take the square root of both sides. x^2=-6+2y

OpenStudy (anonymous):

i assume youre solving for y so subtract 2 form both sides, then divide by -6

OpenStudy (anonymous):

I'm trying to get it to x^2=4py form, so subtracting 2 from both sides wouldn't work because then I would have x^2-2 and that's not what I want. And if I take the square root of both sides, then I lose the x^2 on the left. Right? Or am I doing it wrong?

OpenStudy (anonymous):

i dont know that form

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yeah...I wasn't paying attention in class so now I'm trying to figure it out. It's what I want, according to the book and it got me a correct looking answer.